OpenStudy (anonymous):

Factor: 4xy+10x+2y+5

hartnn (hartnn):

what you can see common to 1st 2 terms ? 4xy+10

OpenStudy (anonymous):

The "x"?

hartnn (hartnn):

and ?

hartnn (hartnn):

how about 2 also ? 4=2*2 10=2*5

OpenStudy (anonymous):

so 2x would be common.

hartnn (hartnn):

exactly, factor it out, what u get ?

OpenStudy (anonymous):

2y+5?

hartnn (hartnn):

thats correct, so you have 2x (2x+5)+1 (2x+5), right ? now factor out (2x+5) what u get ?

hartnn (hartnn):

***2x (2y+5)+1 (2y+5), right ?

hartnn (hartnn):

and factor out 2y+5

OpenStudy (anonymous):

so it would be 2x(2y+5)(2x+5)?

hartnn (hartnn):

no.... wait, i'll draw.

hartnn (hartnn):

 too messy ? :P